List all the factors of 40.
List all the factors of 113.

Answer by jim_thompson5910(35256)

Solved by pluggable solver: Factor any number
40 is NOT a prime number: 40 = 2 * 2 * 2 * 5

Work Shown

40 is divisible by 2: 40 = 20 * 2.
20 is divisible by 2: 20 = 10 * 2.
10 is divisible by 2: 10 = 5 * 2.
5 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.



Solved by pluggable solver: Factor any number
113 is a prime number

Work Shown

113 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.